The equation y=a sin(wt-kx) represents a plane wave travelling in a medium along the x-direction, y being the displacement at the point x at time t. Given that x is in meter and t in second, state the units of k and w. What physical quantity does w/k represent? Justify your answer. State whether the wave is travelling in the negative or positive x-direction
1
Expert's answer
2019-06-17T13:09:20-0400

The unit of k is


"m^{-1}"

The unit of w is


"s^{-1}"

w/k represents phase velocity of the wave.


"\dim \frac{w}{k}=\frac{s^{-1}}{m^{-1}}=\frac{m}{s}=\dim v"

The wave is travelling in the positive x-direction.